# Large Language Models for Software Engineering: A Systematic Literature Review

### v5

@@ 4.3. How do data types influence the selection of data-preprocessing techniques?
Table 7. The data preprocessing procedure for code-based datasets.
Preprocessing techniques Description Examples References
Data extraction Extract relevant code blocks for specific software engineering tasks from code-based datasets, considering different granularities and requirements. Token-level, statement-level, method-level, file-level, or project-level. (Chan et al., 2023) (Ghadhab et al., 2021) (Kang et al., 2023) (Yang et al., 2023c)
Unqualified data deletion Eliminate unqualified data by applying filtering rules to retain appropriate samples, ensuring the dataset’s quality and relevance for various software engineering tasks. Retain only code longer than a certain number of lines, or remove files or methods that contain a certain keyword. (Jana et al., 2023) (Li et al., 2021) (Prenner and Robbes, 2021) (Shi et al., 2022)
Duplicated instance deletion Remove duplicated instances from the dataset to ensure data integrity and prevent redundancy in the training process. Remove near-duplicate code samples using certain deduplication algorithms. (Ciniselli et al., 2021) (Scao et al., 2022) (Xu et al., 2022)
Data compilation Compile the code to get correctly compilable files. Converting java files to .class files throughout the compilation process. (Chan et al., 2023) (Mastropaolo et al., 2022a)
Uncompilable data deletion Remove non-compilable code fragments. Remove code fragments that fail compilation, such as those with syntax errors. (Tian et al., 2023a)
Code representation Represented as tokens. Tokenize source or binary code as tokens. (Xu et al., 2022) (Li et al., 2023d)
Represented as trees. Parses source or binary code into AST. (Al-Kaswan et al., 2023) (Izadi et al., 2022)
Represented as graphs. Generate source or binary code as PDG (CFG, CG). (Ma et al., 2023a) (Zhang et al., 2023c)
Data segmentation Split the dataset for use in a training model, validation model, or test model. Divide the data set according to certain rules, which can be divided into training sets, validation sets, or test sets. (Ciniselli et al., 2021) (Weyssow et al., 2023)
@@
For the training and application of LLMs, the raw dataset needs to be subjected to data processing to obtain a clean and suitable dataset for model training. The data processing steps (Manh et al., 2023; Lee et al., 2022) involve operations such as data cleaning, noise removal, normalization, etc. To ensure consistency and quality of the data, different data types may require different processing methods to improve the performance and effectiveness of LLMs in SE tasks. In this section, we aim to detail the data preprocessing procedures for the two most used types of datasets, i.e., code-based datasets and text-based datasets.
The data preprocessing procedure for code-based datasets. We now summarize the process of preprocessing a code-based dataset, which consists of seven steps. Table 7 describes the individual data processing steps in detail and gives examples. The first step is data extraction, which involves retrieving relevant code segments from different sources such as software repositories or version control systems (Kang et al., 2023; Yang et al., 2023c). Depending on the requirements of the research task (Mastropaolo et al., 2021b; Yuan et al., 2023b), code segments can be extracted at different levels of granularity, ranging from individual methods and functions to entire source code files or even complete software projects. The next step is to remove any code segments that do not meet predefined criteria or quality standards (Li et al., 2021; Shi et al., 2022; Prenner and Robbes, 2021). This filtering process ensures that the extracted code is relevant to the specific SE task under study, thus eliminating incomplete or irrelevant code snippets. To avoid introducing bias and redundancy during model training, the third step involves removing duplicate instances (Zhao et al., 2021; Ciniselli et al., 2021; Xu et al., 2022). Any duplicate code instances are identified and removed from the dataset, thus increasing the diversity and uniqueness of the data. After the data extraction and filtering steps, the fourth step, data compilation, comes into play. The extracted and filtered code segments are merged and compiled into a unified code dataset. This compilation process simplifies data storage and access and facilitates subsequent analysis and model training (Chan et al., 2023; Mastropaolo et al., 2022a). In the fifth step, the problem of invalid or non-executable code is solved by removing data that cannot be compiled. Any code segments that cannot be compiled or executed are removed from the dataset to ensure that the remaining code instances are valid and usable during model training and evaluation. The sixth step is code representation, which consists of converting the code segments into a suitable representation that can be processed by the LLMs. This conversion can take different forms: token-based representation involves tokenizing the source or binary code into distinct tokens; tree-based representation parses the code into Abstract Syntax Trees (AST); and graph-based representation generates a Program Dependence Graph (PDG), encompassing Control Flow Graphs (CFG) and Call Graphs (CG). Finally, in the “data segmentation” step, the preprocessed dataset is partitioned into different subsets for training, validation, and testing (Ciniselli et al., 2021; Weyssow et al., 2023). The training set is used to train the LLM, the validation set helps to tune the hyperparameters and optimize the model performance, and the testing set evaluates the model’s ability on unseen data. By strictly adhering to these seven preprocessing steps, researchers can create structured and standardized code-based datasets, thus facilitating the effective application of LLMs for a variety of SE tasks such as code completion, error detection, and code summarization.
Table 8. The data preprocessing procedure for text-based datasets.
Preprocessing techniques Description Examples References
Data extraction Extract valid text from documentation according to different software engineering tasks. Bug report, requirement documentation, code comments, API documentation, etc. (Yang et al., 2023c) (Ciborowska and Damevski, 2022) (Ezzini et al., 2022) (Ciborowska and Damevski, 2023)
Initial data segmentation Split data into different categories as required. Split data into sentences or words. (He et al., 2023) (Kou et al., 2023a)
Unqualified data deletion Delete invalid text data according to the specified rules. Remove the source code from the bug report. (Gomes et al., 2023) (Mukherjee and Hellendoorn, 2023) (Prenner and Robbes, 2021)
Text preprocessing Preprocessing operations on text. Remove certain symbols and words, or convert all content to lowercase. (Rahmani et al., 2023) (Wang et al., 2020c)
Duplicated instance deletion Remove duplicate samples from the dataset. Utilize the deduplication algorithm from CodeSearchNet (Husain et al., 2019) to remove nearly duplicate methods. (Xu et al., 2022)
Data tokenization Use token-based text representation. Tokenize the texts, sentences, or words into tokens. (Luo et al., 2022)
Data segmentation Split the dataset for use in a training model, validation model, or test model. Divide the data set according to certain rules, which can be divided into training sets, validation sets, or test sets. (Lee et al., 2022)
@@
The data preprocessing procedure for text-based datasets. As displayed in Table 8, the steps of text-based dataset preprocessing consist of a total of seven steps, but there are some differences from the code-based dataset preprocessing steps. The process begins with data extraction (Yang et al., 2023c; Ciborowska and Damevski, 2022; Ezzini et al., 2022; Ciborowska and Damevski, 2023), where relevant text is carefully extracted from SE documentation from a variety of sources, including bug reports (Ciborowska and Damevski, 2023), requirements documents (Kolthoff et al., 2023), code comments (Prenner and Robbes, 2021), and API documentation (Khan et al., 2021). This step ensures that the dataset captures diverse, task-specific textual information. After data extraction, the text is initially segmented and categorized according to the specific requirements of the research task. For example, the text can be segmented into sentences or further broken down into individual words as needed for analysis (He et al., 2023; Kou et al., 2023a). To ensure the quality and relevance of the dataset, substandard data deletion is performed to eliminate any invalid or irrelevant text. For example, the dataset used by Lee et al. (Lee et al., 2022) was constructed from bug reports, and in the “unqualified data deletion” process the researchers filtered out bug reports with fewer than 15 words because the text was too short to contain contextual information. Next, preprocessing operations are performed on the text to standardize and clean it. Common preprocessing steps include removing certain symbols, stop words, and special characters (Rahmani et al., 2023; Wang et al., 2020c). This standardized form of text facilitates the efficient processing of LLMs. To avoid introducing bias and redundancy in the dataset, we eliminated duplicate instances by removing any duplicate text samples (Xu et al., 2022). This step enhances the diversity of the dataset and helps the model to generalize better to new inputs. “Data tokenization” is a key step in preparing the text for LLMs (Luo et al., 2022). Text is labeled into smaller units, such as words or subwords, so that LLMs are easier to manage and process efficiently. Finally, the preprocessed dataset is partitioned into different subsets, usually including a training set, a validation set, and a test set.
Table 9. The various input forms of LLMs proposed in prior studies.
Category Input forms # Studies Total References
Token-based input Code in tokens 64 192 (Ahmed et al., 2023) (Arakelyan et al., 2023) (Bui et al., 2023) (Cassano et al., 2023) (Chan et al., 2023) (Chen et al., 2022) (Chen et al., 2021b) (Chen et al., 2023b) (Chen et al., 2023a) (Ciniselli et al., 2021) (Dou et al., 2023) (Fakhoury et al., 2023) (Fatima et al., 2022) (Geng et al., 2024) (Ghadhab et al., 2021) (Gupta et al., 2023) (Huang et al., 2023d) (Huang et al., 2023c) (Izadi et al., 2022) (Jana et al., 2023) (Jesse et al., 2022) (Kanade et al., 2020) (Kang et al., 2023) (Khanfir et al., 2023) (Lahiri et al., 2022) (Lai et al., 2023) (Lajkó et al., 2022) (Le-Cong et al., 2023) (Li et al., 2021) (Li et al., 2023b) (Liu et al., 2020) (Liu et al., 2023e) (Mastropaolo et al., 2022a) (Mastropaolo et al., 2022b) (Nguyen et al., 2023) (Ochs et al., 2023) (Pan et al., 2023a) (Paul et al., 2023a) (Paul et al., 2023b) (Rao et al., 2023a) (Saberi et al., 2023b) (Schäfer et al., 2023) (Sharma et al., 2022) (Shi et al., 2022) (Shen et al., 2022) (Siddiq et al., 2023b) (Tang et al., 2023b) (Tian et al., 2020) (Tian et al., 2023a) (Tufano et al., 2023) (Wan et al., 2022b) (Widjojo and Treude, 2023) (Xia et al., 2022) (Xia and Zhang, 2023a) (Xie et al., 2023) (Wong et al., 2023) (Yu et al., 2023) (Zeng et al., 2022) (Zhang et al., 2023d) (Zhang et al., 2023g) (Zhang et al., 2023a) (Zhang et al., 2023f) (Zheng et al., 2023a) (Zhu et al., 2023)
Text in tokens 99 (Alhamed and Storer, 2022) (Alqarni and Azim, 2022) (Biswas et al., 2020) (Cassano et al., 2023) (Chen et al., 2023c) (Ciborowska and Damevski, 2022) (Ciborowska and Damevski, 2023) (Deng et al., 2023d) (Deng et al., 2023c) (Dibia et al., 2022) (Dinh et al., 2023) (Döderlein et al., 2022) (Dong et al., 2023) (Du et al., 2023) (Ezzini et al., 2022) (Feng and Chen, 2023) (Fu and Tantithamthavorn, 2022) (Gao et al., 2023) (Gilbert et al., 2023) (Gomes et al., 2023) (He et al., 2023) (He et al., 2022) (Helmeczi et al., 2023) (Hendrycks et al., 2021) (Henkel et al., 2021) (Hey et al., 2020) (Huang et al., 2023a) (Huang et al., 2023b) (Isotani et al., 2021) (Jain et al., 2022) (Jiang et al., 2023b) (Jiang et al., 2023a) (Jin et al., 2023) (Jones and Steinhardt, 2022) (Kang et al., 2022) (Ke et al., 2023) (Khakhar et al., 2023) (Khan et al., 2021) (Khan and Uddin, 2022) (Koide et al., 2023) (Kolthoff et al., 2023) (Kou et al., 2023a) (Kou et al., 2023b) (Kuznia et al., 2022) (Lee et al., 2022) (Li et al., 2022d) (Li et al., 2023c) (Li et al., 2023e) (Lin et al., 2021) (Liu et al., 2023d) (Luitel et al., 2023) (Luo et al., 2022) (Luo et al., 2023) (Mandal et al., 2023) (Mastropaolo et al., 2021a) (Mastropaolo et al., 2021b) (Moharil and Sharma, 2022) (Moharil and Sharma, 2023) (Mukherjee and Hellendoorn, 2023) (Ouyang et al., 2023) (Patil et al., 2023) (Pearce et al., 2021) (Pegolotti et al., 2023) (Qian et al., 2023) (Saieva et al., 2023) (Scao et al., 2022) (Schlag et al., 2023) (Sghaier and Sahraoui, 2023) (Siddiq et al., 2023a) (Singla, 2023) (Sobania et al., 2023) (Sun et al., 2023) (Tan et al., 2023) (Tang et al., 2023a) (Tarassow, 2023) (Thapa et al., 2022) (Tian et al., 2023b) (Tihanyi et al., 2023) (Tu et al., 2023) (Vikram et al., 2023) (Von der Mosel et al., 2022) (Wang et al., 2023b) (Wang et al., 2022b) (Wang et al., 2023e) (Wang et al., 2020c) (Wang et al., 2022c) (Wei et al., 2022a) (Weyssow et al., 2023) (Wu et al., 2023a) (Xu et al., 2023) (Yang et al., 2022c) (Ye et al., 2023) (Yuan et al., 2023b) (Zan et al., 2022a) (Zan et al., 2022b) (Zhang et al., 2022b) (Zhang et al., 2020b) (Zhao et al., 2023a) (Zhou et al., 2023a)
Code and text in tokens 29 (Fan et al., 2023a) (Gu et al., 2022) (Hu et al., 2023) (Li et al., 2022f) (Li et al., 2023d) (Li et al., 2023a) (Liu et al., 2023a) (Liu et al., 2023c) (Manh et al., 2023) (Nijkamp et al., 2023) (Olausson et al., 2023) (Prenner and Robbes, 2021) (Pudari and Ernst, 2023) (Rahmani et al., 2023) (Saberi et al., 2023b) (Sadik et al., 2023) (Sakib et al., 2023) (Salza et al., 2022) (Shirafuji et al., 2023) (Tufano et al., 2022) (Wang et al., 2023c) (Xia and Zhang, 2023b) (Yetiştiren et al., 2023) (Yuan et al., 2022) (Zan et al., 2023a) (Zhang et al., 2022a) (Zhang et al., 2022c) (Zhang et al., 2023e) (Zheng et al., 2023b)
Tree/Graph-based input Code in tree structure 5 7 (Al-Kaswan et al., 2023) (Izadi et al., 2022) (Ochs et al., 2023) (Shi et al., 2022) (Zheng et al., 2023a)
Code in graph structure 2 (Ma et al., 2023a) (Zhang et al., 2023c)
Pixel-based input Pixel 1 1 (Nasir et al., 2023)
Hybrid-based input Hybrid input forms 1 1 (Niu et al., 2022)
@@ 4.4. What input formats are the datasets for LLM training converted to?
Once suitable datasets have been carefully chosen and clean data has been achieved through the preprocessing steps, the next critical aspect is the transformation of the data into appropriate formats that can effectively serve as inputs for LLMs. Table 9 shows four distinct data input types that emerged during the research: Token-based input, Tree/Graph-based input, Pixel-based input, and Hybrid-based input.
Token-based input. Token-based input (Ahmed et al., 2023; Al-Kaswan et al., 2023; Alqarni and Azim, 2022; Arakelyan et al., 2023) involves representing code and text as sequences of tokens, which are smaller units like words or subwords. Code in tokens refers to the representation of code snippets broken down into meaningful tokens, allowing the LLMs to understand programming language syntax and semantics at a fine-grained level. Text in tokens refers to the tokenization of textual data, such as documentation, bug reports, or requirements, enabling the LLMs to process and analyze natural language descriptions effectively. Code and text in tokens combine both code and its associated textual context, allowing the model to capture the relationships between code elements and their descriptions.
Tree/Graph-based input. Tree-based input (Ma et al., 2023a; Ochs et al., 2023; Zhang et al., 2023c) represents code as hierarchical tree structures, capturing the syntactic relationships between code elements. Each node in the tree represents a code element, and the edges represent the hierarchical nesting of control flow statements and other code structures. This form of input allows the LLMs to understand the code’s hierarchical structure and perform tasks like code completion and bug fixing. Graph-based input represents code as a graph structure, where nodes represent code elements and edges represent the relationships between them. Unlike trees, graphs allow more flexible and complex relationships between code elements, enabling the model to capture non-linear dependencies in the code. This form of input is used in tasks like code summarization and vulnerability detection by considering the code’s intricate relationships.
@@
During our investigation of LLM-based models for SE tasks, we observed distinct trends in the usage of different input forms during the training process. Token-based input forms, namely code in tokens and text in tokens were the most prevalent, collectively constituting approximately 95.52% of the studies [^4] This refers to studies that explicitly state input forms of LLMs, i.e., a total of 201 papers as shown in Table 9.. Specifically, code in tokens was widely adopted in 64 studies, accounting for approximately 31.84% of the total studies, demonstrating its popularity as a primary choice for representing code snippets. This approach allowed LLMs to grasp programming language syntax and semantics effectively, making it suitable for a wide range of code-related tasks. Similarly, text in tokens was utilized in 99 studies, comprising around 49.25% of the total studies. This input form allowed LLMs to process natural language descriptions, bug reports, and documentation with greater efficiency and accuracy. The popularity of token-based input forms underscores their significance in leveraging the power of LLMs for software engineering applications.
In contrast, tree/graph-based input forms, such as code in tree-structure, were used in only seven studies, making up approximately 3.48% of the total. Although less prevalent, this input type emerged as a promising choice to represent the hierarchical structure and syntactic relationships within code. Its adoption indicated an ongoing exploration of tree-based representations in specialized tasks, such as code completion and bug fixing.
Pixel-based input and hybrid-based input were relatively less common, each found in one study, contributing approximately 0.5% of the total studies each. While their adoption rates were lower, these input forms presented intriguing possibilities for specific applications. Pixel-based input offered a unique visual representation of code, potentially advantageous for code translation tasks. Meanwhile, hybrid-based input, combining multiple modalities (e.g., code in tree structure and text in tokens in Niu et al.’s work (Niu et al., 2022)), showcased the potential for enhancing code comprehension tasks by offering diverse perspectives for the models to learn from.
@@
RQ2 - Summary (1) We divided the datasets into four categories based on the source of data: open-source, collected, constructed, and industrial datasets. The use of open-source datasets is the most prevalent, constituting approximately 59.35% of the 214 papers that explicitly state the dataset. (2) We categorized the data types within all datasets into five groups: code-based, text-based, graph-based, software repository-based, and combined. Text-based and code-based types are the most frequently used in applying LLMs to SE tasks. This pattern indicates that LLMs are particularly adept at handling text and code-based data in SE tasks, leveraging their natural language processing capabilities. (3) We summarized the data preprocessing procedures for different data types and found several common preprocessing procedures, i.e., data extraction, unqualified data deletion, duplicated instance deletion, and data segmentation.

### v6

@@ 4.3. How do data types influence the selection of data-preprocessing techniques?
For the training and application of LLMs, the raw dataset needs to be subjected to data processing to obtain a clean and suitable dataset for model training. The data processing steps (Manh et al., 2023; Lee et al., 2022) involve operations such as data cleaning, noise removal, normalization, etc. To ensure consistency and quality of the data, different data types may require different processing methods to improve the performance and effectiveness of LLMs in SE tasks. In this section, we aim to detail the data preprocessing procedures for the two most used types of datasets, i.e., text-based datasets and code-based datasets.
Figure 7. The data preprocessing procedure for text-based datasets.
Figure 8. The data preprocessing procedure for code-based datasets.
@@
The data preprocessing procedure for text-based datasets. As displayed in Fig.  7, the steps of text-based dataset preprocessing consist of seven steps in total, yet there are some differences from the code-based dataset preprocessing steps. The process begins with data extraction (Yang et al., 2023f; Ciborowska and Damevski, 2022; Ezzini et al., 2022; Ciborowska and Damevski, 2023), where relevant text is carefully extracted from SE documentation from a variety of sources, including bug reports (Ciborowska and Damevski, 2023), requirements documents (Kolthoff et al., 2023), code comments (Prenner and Robbes, 2021), and API documentation (Khan et al., 2021). This step ensures that the dataset captures diverse, task-specific textual information. After data extraction, the text is initially segmented and categorized according to the specific requirements of the research task. For example, the text can be segmented into sentences or further broken down into individual words as needed for analysis (He et al., 2023; Kou et al., 2023a). To ensure the quality and relevance of the dataset, substandard data deletion is performed to eliminate any invalid or irrelevant text. For example, the dataset used by Lee et al. (Lee et al., 2022) was constructed from bug reports, and in the “unqualified data deletion” process the researchers filtered out bug reports with fewer than 15 words because the text was too short to contain contextual information. Next, preprocessing operations are performed on the text to standardize and clean it. Common preprocessing steps include removing certain symbols, stop words, and special characters (Rahmani et al., 2023; Wang et al., 2020c). This standardized form of text facilitates the efficient processing of LLMs. To avoid introducing bias and redundancy in the dataset, researchers eliminated duplicate instances by removing any duplicate text samples (He et al., 2023; Kou et al., 2023a; Xu et al., 2022). This step enhances the diversity of the dataset and helps the model to generalize better to new inputs. “Data tokenization” is a key step in preparing the text for LLMs (Luo et al., 2022). Text is labeled into smaller units, such as words or subwords, so that LLMs are easier to manage and process efficiently. Finally, the preprocessed dataset is partitioned into different subsets, usually including a training set, a validation set, and a test set.
The data preprocessing procedure for code-based datasets. We now summarize the process of preprocessing a code-based dataset, which consists of seven steps. Fig.  8 describes the individual data processing steps in detail and gives examples. The first step is data extraction, which involves retrieving relevant code segments from different sources such as software repositories or version control systems (Kang et al., 2023b; Yang et al., 2023f). Depending on the requirements of the research task (Mastropaolo et al., 2021b; Yuan et al., 2023b), code segments can be extracted at different levels of granularity, ranging from individual methods and functions to entire source code files or even complete software projects. The next step is to remove any code segments that do not meet predefined criteria or quality standards (Li et al., 2021; Shi et al., 2022; Prenner and Robbes, 2021). This filtering process ensures that the extracted code is relevant to the specific SE task under study, thus eliminating incomplete or irrelevant code snippets. To avoid introducing bias and redundancy during model training, the third step involves removing duplicate instances (Zhao et al., 2021; Ciniselli et al., 2021; Xu et al., 2022). Any duplicate code instances are identified and removed from the dataset, thus increasing the diversity and uniqueness of the data. After the data extraction and filtering steps, the fourth step, data compilation, comes into play. The extracted and filtered code segments are merged and compiled into a unified code dataset. This compilation process simplifies data storage and access and facilitates subsequent analysis and model training (Chan et al., 2023; Mastropaolo et al., 2022a). In the fifth step, the problem of invalid or non-executable code is solved by removing data that cannot be compiled. Any code segments that cannot be compiled or executed are removed from the dataset to ensure that the remaining code instances are valid and usable during model training and evaluation. The sixth step is code representation, which consists of converting the code segments into a suitable representation that can be processed by the LLMs. This conversion can take different forms: token-based representation involves tokenizing the source or binary code into distinct tokens; tree-based representation parses the code into Abstract Syntax Trees (AST); and graph-based representation generates a Program Dependence Graph (PDG), encompassing Control Flow Graphs (CFG) and Call Graphs (CG). Finally, in the “data segmentation” step, the preprocessed dataset is partitioned into different subsets for training, validation, and testing (Ciniselli et al., 2021; Weyssow et al., 2023b). The training set is used to train the LLM, the validation set helps to tune the hyperparameters and optimize the model performance, and the testing set evaluates the model’s ability on unseen data. By strictly adhering to these seven preprocessing steps, researchers can create structured and standardized code-based datasets, thus facilitating the effective application of LLMs for a variety of SE tasks such as code completion, error detection, and code summarization.
It is worth emphasizing that the order of these steps is not fixed and can be adjusted based on the specific research task and its associated requirements. Researchers need to carefully consider the objectives, characteristics of the dataset, and the desired outcomes when determining the optimal sequence for these preprocessing techniques.
@@ 4.4. What input formats are the datasets for LLM training converted to?
Once suitable datasets have been carefully chosen and clean data has been achieved through the preprocessing steps, the next critical aspect is the transformation of the data into appropriate formats that can effectively serve as inputs for LLMs. Table 8 shows four distinct data input types that emerged during the research: Token-based input, Tree/Graph-based input, Pixel-based input, and Hybrid-based input. We now detail each as follows:
Table 8. The various input forms of LLMs proposed in prior studies. See Appendix B for the full table including references.
Category Input forms Total
Token-based input Text in tokens (150) Code in tokens (118) 347
Code and text in tokens (78)
Tree/Graph-based input Code in tree structure (2) Code in graph structure (3) 5
Pixel-based input Pixel (1) 1
Hybrid-based input Hybrid input forms (2) 2
Token-based input. Token-based input (Ahmed et al., 2024; Al-Kaswan et al., 2023; Arakelyan et al., 2023) involves representing code and text as sequences of tokens, which are smaller units like words or subwords. Text in tokens refers to the tokenization of textual data, such as documentation, bug reports, or requirements, enabling the LLMs to process and analyze natural language descriptions effectively. Code and text in tokens combine both code and its associated textual context, allowing the model to capture the relationships between code elements and their descriptions. Code in tokens refers to the representation of code snippets broken down into meaningful tokens, allowing the LLMs to understand programming language syntax and semantics at a fine-grained level.
Tree/Graph-based input. Tree-based input (Ma et al., 2023a; Ochs et al., 2023; Zhang et al., 2023j) represents code as hierarchical tree structures, capturing the syntactic relationships between code elements. Each node in the tree represents a code element, and the edges represent the hierarchical nesting of control flow statements and other code structures. This form of input allows the LLMs to understand the code’s hierarchical structure and perform tasks like code completion and bug fixing. Graph-based input represents code as a graph structure, where nodes represent code elements and edges represent the relationships between them. Unlike trees, graphs allow more flexible and complex relationships between code elements, enabling the model to capture non-linear dependencies in the code. This form of input is used in tasks like code summarization and vulnerability detection by considering the code’s intricate relationships.
@@
During our investigation of LLM-based models for SE tasks, we observed distinct trends in the usage of different input forms during the training process. Token-based input forms, namely code in tokens and text in tokens were the most prevalent, collectively constituting approximately 97.75% of the studies [^5] This refers to studies that explicitly state input forms of LLMs, i.e., a total of 355 papers as shown in Table 8.. Specifically, code in tokens was widely adopted in 118 studies, accounting for approximately 33.24% of the total studies, demonstrating its popularity as a primary choice for representing code snippets. This approach allowed LLMs to grasp programming language syntax and semantics effectively, making it suitable for a wide range of code-related tasks. Similarly, text in tokens was utilized in 150 studies, comprising around 42.25% of the total studies. This input form allowed LLMs to process natural language descriptions, bug reports, and documentation with greater efficiency and accuracy. The popularity of token-based input forms underscores their significance in leveraging the power of LLMs for software engineering applications.
In contrast, tree/graph-based input forms, such as code in tree-structure, were used in only seven studies, making up approximately 1.4% of the total. Although less prevalent, this input type emerged as a promising choice to represent the hierarchical structure and syntactic relationships within code. Its adoption indicated an ongoing exploration of tree-based representations in specialized tasks, such as code completion and bug fixing.
Pixel-based input and hybrid-based input were relatively less common, each found in one study, contributing approximately 0.28% of the total studies each. While their adoption rates were lower, these input forms presented intriguing possibilities for specific applications. Pixel-based input offered a unique visual representation of code, potentially advantageous for code translation tasks. Meanwhile, hybrid-based input, combining multiple modalities (e.g., code in tree structure and text in tokens in Niu et al.’s work (Niu et al., 2022)), showcased the potential for enhancing code comprehension tasks by offering diverse perspectives for the models to learn from.
@@
RQ2 - Summary
(1) We divided the datasets into four categories based on the source of data: open-source, collected, constructed, and industrial datasets. The use of open-source datasets is the most prevalent, constituting approximately 62.83% of the 374 papers that explicitly state the dataset. (2) We categorized the data types within all datasets into five groups: code-based, text-based, graph-based, software repository-based, and combined. Text-based and code-based types are the most frequently used in applying LLMs to SE tasks. This pattern indicates that LLMs are particularly adept at handling text and code-based data in SE tasks, leveraging their natural language processing capabilities. (3) We summarized the data preprocessing procedures for different data types and found several common preprocessing procedures, i.e., data extraction, unqualified data deletion, duplicated instance deletion, and data segmentation.